(1 point) Find the QR factorization of 4 -14 6 -14 12 -21 M = ME
Answer #1

Given mat vix is - M =[me m2] 4 -14 6 -14 12 -21 Let M = QR - 4 -14 6 -14 12 -21 [9; 12 TR1 R12 R22 o o the R First column ofใน =1-11 2/3 - 14 -(-2| ปา -21 6/7 ใว้ -19 +2 - 19 +12 -2f+29 % -6 2 3 ใน - 1 ใน R22 Now R22 || 2 || =? ใน 2 3 ใน - 61) -25 วThas QR factorisation of M is 나 -1, - M- 27 19 228 - 7 매 7 뼡 12 -21 6/7 37

Please thumbs up ?